OpenStudy (amtran_bus):

Since current is v/r is you had a 50 ohm and 10 ohm resistor, which would have most current? Show work.

OpenStudy (amtran_bus):

I know it is a physics problem but it requires the math and I need the steps

OpenStudy (amtran_bus):

@Jonask @jhonyy9 @robtobey

OpenStudy (amtran_bus):

v is voltage and r is resistance

OpenStudy (anonymous):

you can pick any value for v and use that value for calculating the current in the two resistors...

OpenStudy (amtran_bus):

Well, if it was a 9V battery I=9/10=.9 I=9/50=.18 But I thought the more resistance the less current...

OpenStudy (amtran_bus):

@kausarsalley

OpenStudy (anonymous):

it depends how you connect them in 1 series 2 parrallel 3 seperate circuit with same voltage

OpenStudy (amtran_bus):

This is in parallel

OpenStudy (anonymous):

case 1 they have the same current case 2 smaller resistor has more current since we are using \[\frac{ 1 }{ R_p }=\frac{ 1 }{ R_1 }+\frac{ 1 }{ R_2 }\] current is split accorind to their ratio jus imagine a traffic with lots of cars(resistance) and a one without current takes the easy way out ,smaller traffic(resistanvce)

OpenStudy (amtran_bus):

The attatchment has the problem and the key said B is that right

OpenStudy (anonymous):

yes....the more the resistance, the lesser the current

OpenStudy (anonymous):

yes

OpenStudy (amtran_bus):

I=9/10=.9 I=9/50=.18

OpenStudy (anonymous):

just as demonstrated in your calculation....

OpenStudy (amtran_bus):

THanks!

OpenStudy (anonymous):

do you understand it now???